Circular No. 6042

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994W IN NGC 4041 G. Cortini and M. Villi, Valbura, Italy, report their visual discovery

on July 29.85 UT of an apparent supernova of mag ∼ 13.5 located ∼ 1′ north of the center of NGC 4041 (α = 11^{h}59^{m}.6, δ = +62^{o}25′, equinox 1950.0). The object was fainter than m_{v} ∼ 14.5 (not seen) on July 11.8. R. Barbon and U. Munari, Asiago Astrophysical Observatory; and A. Bragaglia, Bologna Astronomical Observatory (BAO), report that CCD images secured on July 30.9 with the 1.5-m telescope of BAO show SN 1994W to be superimposed on a galaxy arm and located 18′′ north, 9′′ west of the galaxy’s nucleus. C. Pollas and Liu Zongli also confirm the existence of this new object on panchromatic films obtained July 30 with the Schmidt telescope at Observatoire de la Côte d’Azur.

JUPITER AND COMET 1993e M. J. Klein and S. Gulkis, Jet Propulsion Laboratory, report a pro-

nounced increase in the microwave radio emission from Jupiter apparently associated with the bombardment of the planet by P/Shoemaker-Levy 9: “The increase in radio intensity began soon after July 16 and peaked July 23, when the flux density reached 7.5 Jy. This measurement represents a 25-percent increase over the baseline measurements that began in 1993 Dec. and continued until 1994 July 16.2 UT, just hours before the impact of fragment A. The flux density has been slowly fading each day since July 23, but is still much higher than normal on July 27. The flux density reported here is not the peak flux; it is the A_{0} term of the Fourier-series fit (1989, NASA SP-494, pp. 151–155) to the System-III rotation (beaming) curve repeatedly measured with the Goldstone 70-m antenna during the baseline measurements. The current observations are being made with a 34-m antenna at NASA’s Goldstone Deep Space Communication Complex. The prime frequency is 2295 MHz with right circular polarized feeds. Simultaneous measurements are being made at 8420 MHz (left circular polarized), but data have not yet been reduced. Observations will continue for at least several weeks. These new results are consistent with, and tend to corroborate, the 50-percent increase that was reported at 843 MHz by Hunstead and Campbell-Wilson (IAUC 6038).”

1994 July 30 Daniel W. E. Green